OpenStudy (anonymous):

If the amount of air in a spherical balloon is 288 pi cu. in., what is the radius of the balloon?

OpenStudy (anonymous):

6 inches 7.3 inches 72 inches

OpenStudy (mertsj):

\[V=\frac{4}{3}\pi r^3\]

OpenStudy (mertsj):

Replace V with 288 pi and solve for r

OpenStudy (anonymous):

how

OpenStudy (mertsj):

Multiply both sides by 3/4 and tell me what you get.

OpenStudy (anonymous):

288*3/4 = 216

OpenStudy (mertsj):

\[216\pi =\pi r^3\]

OpenStudy (mertsj):

That's what you should have after you multiply both sides by 3/4

OpenStudy (mertsj):

Now divide both sides by pi

OpenStudy (mertsj):

What do you get?

OpenStudy (anonymous):

63.5

OpenStudy (mertsj):